OpenStudy (anonymous):

Simplify radical? -√324?

OpenStudy (anonymous):

i18 where i=imaginary number.

OpenStudy (lgbasallote):

not exactly...

OpenStudy (lgbasallote):

\[\sqrt{-324} = 18i\] HOWEVER \[-\sqrt{324} \implies -(\sqrt{324}) \implies -(18)\] make sense?

OpenStudy (anonymous):

thanks!!!

OpenStudy (anonymous):

Ah! I can't believe I forgot about my negative, nice catch igbasallote!!

OpenStudy (lgbasallote):

haha ^_^